# and quantum information science

### v1

@@ 5 Paradigms for complexity I: Nielsen complexity
As mentioned in the previous section, one can define two notions of complexity for a unitary operator: exact and approximate. Each notion has a downside. For starters, exact complexity can behave counterintuitively. For example, a unitary close to the identity can have a large complexity as obtaining it exactly might require fine tuned combination of very many gates. Meanwhile, approximate complexity involves a tolerance, which may seem arbitrary. One can resolve both issues by defining complexity in terms of a continuous quantity.
Nielsen et al. identified the optimal implementation of a global unitary $U$, using complexity geometry [78, 79, 80]. Complexity geometry is a unitary-group manifold on whose tangent space a norm is defined. A continuous trajectory on the manifold results from evolution under a time-dependent Hamiltonian. The manifold’s origin represents the identity operator, and another point represents $U$. A geodesic connects the points. In terms of the geodesic’s length, Nielsen et al. lower- and upper-bounded discrete quantum circuits’ complexities [10]. However, one can define the geodesic’s length as a stand-alone notion of the unitary’s complexity: Nielsen’s complexity. Nielsen’s complexity has several appealing features:
@@
    In Nielsen’s framework, evolution along a continuous trajectory decomposes into a sequence of infinitesimal steps. This decomposition is Trotterization, a technique for simulating evolution under a time-independent Hamiltonian, using a circuit (using evolution under a time-varying Hamiltonian) [81]. Hence Nielsen’s complexity is related to an important quantum-computational concept.
@@
    Using Nielsen’s geometry, one can upper-bound the approximate circuit complexity, and lower-bound the exact circuit complexity, discussed in Sec. 4.
@@
    Nielsen’s complexity is a geometric object. Therefore, one can apply to it Riemannian-geometry (or Finsler-geometry) tools.
@@
    Nielsen’s complexity connects naturally to Hamiltonian control problems, extensively studied in the quantum computation literature.
@@
    One has substantial freedom in equipping Nielsen’s complexity geometry with a distance measure on the unitary manifold. This freedom mimics the choice of a gate set in the exact and approximate circuit complexities’ definitions. A priori, computing Nielsen’s complexity is difficult because complexity geometry is high-dimensional and highly curved. Yet different choices of complexity geometry form equivalence classes whose elements behave the same way at long distances. Within an equivalence class, one can often find members that are only modestly curved [80, 82, 14]. This result facilitates the calculation of Nielsen’s complexity.
@@
    The problem of calculating Nielsen’s complexity can become exactly or approximately soluble when considering unitary representations of certain symmetry groups. Examples include unitary representations of symplectic and orthogonal groups (relevant respectively for free bosonic and fermion quantum systems), the global conformal group, and circuits constructed from the Virasoro algebra generators (relevant for spacetime transformations in CFTs).
@@
    Nielsen’s complexity, being continuous, resonates with continuous features of QFT and CFT and the cMERA tensor network [13, 83] and natural physical systems’ time evolution. The cMERA parallel motivated the earliest work on complexity in free QFT [84, 85].
We start in Sec. 5.1 with the definition of Nielsen’s complexity. Section 5.2 discusses possible choices for a metric on the complexity geometry, and their basic properties. In Sec. 5.3, we discuss the required properties a complexity geometry should have in order to reproduce the linear time dependence and switchback effect for system evolution with a chaotic Hamiltonian. In Sec. 5.4, we discuss a choice of norm on the unitary manifold relevant for distributed computing, which highlights a connection between complexity and the entanglement entropy. We then turn to the studies of Nielsen complexity in QFT in Sec. 5.5.
@@ 5.1 Definition of Nielsen complexity
One can define Nielsen’s complexity on a manifold formed by any Lie group $G$. For simplicity and clarity, we now focus on the special unitary group $\mathrm{SU}(N)$. For the $K$-qubit system discussed in the Sec. 4.1, $N=2^{K}$. However, one can define complexity geometries of every symmetry group. Example symmetry groups include global and local conformal groups relevant to CFT problems, symplectic groups relevant to free bosons, and orthogonal groups relevant to free fermions.
We aim to identify the optimal means of generating a target unitary $U$, starting from the identity operator $\mathbbm{1}$. One generates $U$ via evolution in time, $t$, under a Hamiltonian $H(t)$ that may vary. It is important to stress that this auxiliary time $t$ is a priori distinct from the physical time governing unitary time evolution of a system in question. Let $\{T_{I}\}$ denote a basis for the space of Hermitian operators that form the unitary algebra $\mathfrak{su}(N)$. One can expand the Hamiltonian in terms of the basis elements:
@@
The $Y^{I}(t)$ denote velocities, or control functions. They are the components of the vector that is tangent, at the point $t$, to the trajectory generated by $H(t)$ through the group manifold. Different trajectories represent different ways of generating $U$. We have parameterized each trajectory with $t\in[0,1]$, such that $U(t{=}0)=\mathbbm{1}$ and $U(t{=}1)=U$. At each point along the path, the unitary has the form
@@
The path ordering $\mathcal{\overset{\leftarrow}{P}}$ constructs the trajectory by multiplying the exponential terms from right to left. Denote by $F[Y^{I}(t)]$ the cost function, a norm on the tangent space. Nielsen’s unitary complexity is defined as the minimal distance induced by this norm:
@@
The path $U(t)$ obeys the aforementioned boundary conditions. The cost function encodes the difficulty of extending the trajectory in any given direction (in the tangent space at any given manifold point).
The previous definitions refer purely to the space of unitaries. Next, we define complexity in terms of the Hilbert space of whole-system quantum states. Assume that the $\mathrm{SU}(N)$ elements act on an $N$-dimensional Hilbert space $\mathcal{H}$. We would like to define quantum state complexity that captures the difficulty of producing an arbitrary target state $\lvert\psi_{\rm T}\rangle$ from a fixed reference state $\lvert\psi_{\rm T}\rangle$. We define Nielsen’s state complexity as the least complexity of any special unitary that evolves the reference to the target:
@@
The state complexity is a property of two states, yet the definition invokes the unitary complexity (19). Instead, one might prefer to rewrite the above definition in terms of a metric induced on the Hilbert space, which directly gives a notion of infinitesimal distance between states. To define such a metric, recall that the space of rays is a collection of equivalence classes. Each class consists of the vectors that differ by a global phase. This concept is captured by the stabilizer of a state $\lvert\psi\rangle$. The stabilizer of any state $\lvert\psi\rangle$ is the maximal proper subgroup $\mathrm{SU}(N-1)\times\mathrm{U}(1)$ of $\mathrm{SU}(N)$. The stabilizer is spanned by the generators that perform no physically meaningful operations on $\psi$: they do not affect the physical state, except by changing the global phase. Using the stabilizer, we define equivalence classes as follows. Let $V$ denote any element in the $\lvert\psi\rangle$ stabilizer. Consider right-multiplying any element $U\in\mathrm{SU}(N)$ by $V$. The resulting unitary, $U^{\prime}\coloneqq UV$, lies in the same equivalence class as $U$:
@@
Equation (21) defines a quotient map from the unitary group to the complex projective space $\mathbb{CP}^{N-1}$:
@@
The state complexity (20) induces a norm on the state space, a minimum over the trajectories consistent with the boundary conditions. One can perform such a minimization in each infinitesimal step of a trajectory, to obtain a norm on the space of states, as follows. One fixes $|\psi(t)\rangle$ and its time derivative $|\dot{\psi}(t)\rangle$, then minimizes the cost function over the degrees of freedom in the Hamiltonian $H(t)$. This process induces the norm
$F^{\rm state}[\lvert\psi(t)\rangle,|\dot{\psi}(t)\rangle]\coloneqq\min_{\{\mathrm{stab}\bm{(}\lvert\psi(t)\rangle\bm{)}\}}F[Y^{I}(t)]\,.$ (23)
@@
$\mathcal{C}^{\rm state}_{F}[\lvert\psi_{\rm T}\rangle,\lvert\psi_{\rm R}\rangle]=\min_{{\{|\psi(t)\rangle\,:\,|\psi(0)\rangle=\lvert\psi_{\rm R}\rangle,\,|\psi(1)\rangle=\lvert\psi_{\rm T}\rangle\}}}\int_{0}^{1}dt\;F^{\rm state}[\lvert\psi(t)\rangle,|\dot{\psi}(t)\rangle]\,.$ (24)
Using this formula, we can study geodesics in the state space without referring to unitaries.
@@ 5.2 Geometric features of Nielsen complexity
We discuss the parameter $p$ in the next paragraph. The penalty factors $q_{I}\geq 0$ participate in a one-to-one correspondence with the generators $T_{I}$ of the $\mathfrak{su}(N)$ algebra. The penalty factors quantify the difficulties of moving along certain directions of the tangent space. One incorporates features of discrete circuits into the complexity geometry: one chooses penalty factors that reflect the difficulties of performing certain gates. For instance, few-qubit operators are naturally easier to implement than many-qubit operators. Therefore, it is natural to assign smaller penalty factors to the generators (called easy) that implement few-qubit operations and larger penalties to the generators (called hard) that act on several qubits. The terms easy and hard refer to the magnitude of the penalty factor associated with each generator. A trajectory in the group manifold can be approximated by a circuit whose gates act for short time intervals.
@@
Returning to the cost function (25), we analyze the parameter $p=1$. When $p=1$, $F_{p{=}1,\vec{q}}[Y^{I}]$ has a natural physical interpretation: it counts (with an appropriate measure) the number of gates used to build the continuous trajectory in the group manifold SU($N$). However, $F_{p{=}1,\vec{q}}[Y^{I}]$ has the disadvantage of not being smooth. One therefore cannot apply the calculus of variations to the geodesics. When $p=2$, the norm (25) induces a Riemannian metric on the group manifold, simplifying the study of the geodesics. In this case, the map (22) is a Riemannian submersion [86] (defined in [87]). Heuristically, a Riemannian submersion is a smooth mapping from a higher-dimensional manifold to a lower-dimensional one. This mapping preserves distances in the directions perpendicular to the fibers. [^11] The fibers are points, in the higher-dimensional space, that are mapped to the same point in the lower-dimensional manifold. In our setting, the fibers correspond to the maximal subgroup $\mathrm{SU}(N-1)\times\mathrm{U}(1)$. The submersion maps $\mathrm{SU}(N)$ to $\mathbb{CP}^{N-1}$. These facts provide a systematic means of determining the metric, induced by the cost-function minimization (23), on the state space. The Nielsen-complexity literature has mainly focused on $p=1,2$ [84, 85, 88, 89, 90, 91, 92, 37, 93, 94, 44, 30, 95, 96, 86, 32, 97, 82, 98, 14, 34, 99].
@@
One may wonder whether Nielsen’s complexity is related to the discrete gate complexity. The answer is yes: Nielsen’s complexity provides both upper and lower bounds on the gate complexity, provided the cost function satisfies certain conditions [78, 79]. The authors of [78, 79] further show that, by penalizing enough the generators acting on several qubits, one can build a trajectory that reaches a unitary in the complexity geometry with arbitrary precision. This trajectory can be built by only using generators acting on one or two qubits. The penalties required for this construction typically make the manifold highly curved. However, as we will subsequently review, Nielsen’s complexity is universal at long distances within a certain equivalence class of metrics, and this equivalence class can be shown to also contain moderately-curved metrics. A priori, this fact makes the computation of Nielsen’s complexity easier by using tools from differential geometry. When the motion in the unitary manifold is restricted to a small subgroup, exact or semi-analytic results can be obtained, which we discuss in later subsections.
@@
There is a substantial freedom in choosing a cost function. Are there any common properties shared by different cost functions? A large universality class of cost functions exists: many metrics on the group manifold share the features described below [82, 14]. This class’s constituents may differ significantly at short distances. However, they differ at large distances only polynomially in the path length and in the physical system’s size. To simplify the rest of this subsection, we denote the penalty factors by $q_{k}$ and assume that they depend only on the corresponding gate’s weight, $k$ [^12] A generator has weight $k$ if it acts nontrivially on exactly $k$ qubits. $q_{k}$ should not be confused with the $q_{I}$ introduced in Eq. (25). The notation $q_{k}$ alludes to how all the generators $T_{I}$ that have the same weight $k$ share the same penalty factor. Consider increasing any penalty factor while keeping the others constant. Above a critical value $\bar{q}_{k}$, the complexity ceases to depend on the penalty factor. This happens because one can implement a given unitary operator with weight $k$ via circuits formed from other, cheaper gates. (The reason is as follows. Consider any fixed $m\geq 2$. Using the $m$-local gates, one can reconstruct all the $m^{\prime}$-local gates for every $m^{\prime}$.) Consequently, one can identify a critical schedule, a set $\{\bar{q}_{k}\}$ of the penalty factors’ critical values. The universality class is the set of all the schedules such that $q_{2}=\bar{q}_{2}$ and, for all $m>2$, $q_{m}\geq\bar{q}_{m}$. In the universality class, the metrics may differ greatly at short separations but approximately equal each other at long separations. That is, short and long distance scales decouple. At the same time, the critical schedule is conjectured to be the only universality-class member that behaves uniformly across all the distance scales [82]. In other words, the conjugate points are pushed to infinity, and geodesics are straight lines for an exponential time in the number of qubits. Brown argued that the same universality class consists of metrics whose associated Nielsen complexity (measured by the geodesic length from the origin) is polynomially equivalent to the gate complexity. At the same time, one can always pick a metric inside each universality class such that the unitary manifold, equipped with such metric, has modest curvature [14]. Therefore, one can (in principle) more easily compute the Nielsen complexity using tools from differential geometry.

### v2

@@ 5 Paradigms for complexity I: Nielsen complexity
As discussed in the previous section, one can define two notions of complexity for a unitary operator: exact and approximate. Each notion has a downside. For starters, exact complexity can behave counterintuitively. For example, a unitary close to the identity can have a large complexity as obtaining it exactly might require fine tuned combination of very many gates. On the other hand, approximate complexity involves a tolerance, which may seem arbitrary. One can resolve both issues by defining complexity in terms of a continuous quantity.
[…]
Nielsen complexity has several appealing features:
@@
    In Nielsen’s framework, evolution along a continuous trajectory decomposes into a sequence of infinitesimal steps. This decomposition is equivalent to Trotterization, a technique for building a discrete circuit to simulate Hamiltonian time evolution [96]. Hence Nielsen’s complexity is related to an important quantum-computational concept.
@@
    Using Nielsen’s geometry, one can upper-bound the approximate circuit complexity, and lower-bound the exact circuit complexity, concepts discussed in Sec. 4.
@@
    Nielsen complexity is a geometric object and thus we can try to compute and analyze it with the methods of differential and algebraic geometry.
@@
    Nielsen complexity connects naturally to Hamiltonian control problems that are extensively studied in the quantum computation literature.
@@
    One has substantial freedom in equipping Nielsen’s complexity geometry with a distance measure on the unitary manifold. This freedom mimics the choice of a gate set in the definition of exact and approximate circuit complexity. A priori, it is difficult to compute Nielsen’s complexity because the complexity geometry in interesting cases is generally high-dimensional and highly curved. But different choices of metric can belong to the same equivalence class with the same long-distance behavior. Within an equivalence class, we can sometimes find modestly curved members [95, 97, 19], facilitating the calculation of Nielsen complexity.
@@
    The problem of calculating Nielsen complexity can become exactly or approximately solvable when we consider unitary representations of certain symmetry groups. Examples include unitary representations of symplectic and orthogonal groups (relevant respectively for free bosonic and fermionic quantum systems), the global conformal group, and circuits constructed from the Virasoro algebra of generators (relevant for spacetime transformations in CFTs).
@@
    Nielsen complexity, as a continuous quantity, naturally relates to continuous features of QFT and cMERA tensor networks [18, 98], as well as to physical time evolution. In fact, the parallel with cMERA motivated the earliest works on complexity in free QFT [99, 100].
We start in Sec. 5.1 with the definition of Nielsen complexity. Section 5.2 discusses possible choices of complexity metrics. In Sec. 5.3, we discuss the required properties a complexity geometry should have in order to reproduce the linear time dependence and switchback effect for system evolution with a chaotic Hamiltonian. Section 5.4 presents general results on the time evolution of Nielsen’s complexity on the unitary manifold. In Sec. 5.5, we discuss a choice of norm on the unitary manifold relevant for distributed computing, which highlights a connection between complexity and the entanglement entropy. We then turn to studies of Nielsen complexity in QFT in Sec. 5.6.
@@ 5.1 Definition of Nielsen complexity
We will first define a complexity geometry for the special unitary group $\mathrm{SU}(N)$ with $N=2^{K}$ describing transformations acting on states of the $K$-qubit system in Sec. 4.1. As we will later see, we can also define complexity geometries for the global and local conformal groups relevant to conformal field theories, and for the symplectic and orthogonal groups relevant to free bosons and fermions.
We want to identify the optimal trajectory in $\mathrm{SU}(N)$ for generating a target unitary $U$ from the identity via the action of a Hamiltonian $H(t)$. If $H$ and $t$ are the physical Hamiltonian and time, we can think of this as describing physical time evolution as a trajectory on the unitary group. But $H(t)$ could also be an alternate family of operations carried out over an auxiliary time, perhaps giving a quicker way to reach the target $U$ from the identity. Let $\{T_{I}\}$ be a Hermitian basis for the Lie algebra of the group, $\mathfrak{su}(N)$. We can expand the Hamiltonian in this basis as:
@@
The $Y^{I}(t)$ denote velocities or control functions. They are the components of the vector that is tangent at time $t$ to the trajectory generated by $H(t)$ through the group manifold. Different trajectories represent different ways of generating $U$. We have parameterized each trajectory by $t\in[0,1]$, such that $U(t{=}0)=\mathbbm{1}$ and $U(t{=}1)=U$.
At each point along the path, the generated unitary transformation has the form
@@
The path ordering $\mathcal{\overset{\leftarrow}{P}}$ constructs the trajectory by multiplying the exponential terms from right to left. Let $F[Y^{I}(t)]$ be a cost or distance function defined on infinitesimal displacements on the group manifold via an inner product on the tangent space. The cost function encodes the difficulty of carrying out an operation that displaces the unitary operator in any given direction along the tangent space at a given point in the group manifold. The integrated cost/length of a trajectory is an integral of these infinitesimal costs. Nielsen’s complexity is defined as the minimal cost/length induced by the inner product, evaluated over all trajectories between $\mathbbm{1}$ and $U$:
@@
The above definition of the complexity of unitary transformations induces a notion for the complexity of states in the Hilbert space on which the unitaries act. To construct this quantity, assume that the $\mathrm{SU}(N)$ elements act on an $N$-dimensional Hilbert space $\mathcal{H}$. We would like to define a quantum state complexity that captures the difficulty of producing an arbitrary target state $\lvert\psi_{\rm T}\rangle$ from a fixed reference state $\lvert\psi_{\rm R}\rangle$. Let us define Nielsen state complexity as the lowest complexity of any special unitary transformation that evolves the reference to the target:
@@
It would be helpful if this quantity could be written directly in terms of a metric on the Hilbert space. We will define such a metric below.
To define a metric on the Hilbert space, recall first that quantum states should be thought of as equivalence classes of elements that differ by a global phase. So we should first ask, what transformations in $\mathrm{SU}(N)$ leave a particular state invariant up to a phase, as such transformations should induce zero cost. The set of such transformations is the stabilizer of the state. Think of the given state as a basis element of the Hilbert space. Then transformations that only affect the remaining $N-1$ basis elements leave it invariant. Including multiplication by a phase, the stabilizer of a state $\lvert\psi\rangle$ is therefore the maximal proper subgroup $\mathrm{SU}(N-1)\times\mathrm{U}(1)$ of $\mathrm{SU}(N)$.
We can define equivalence classes in terms of the stabilizer as follows. Let $V$ denote an element in the $\lvert\psi\rangle$ stabilizer. Consider right-multiplying any element $U\in\mathrm{SU}(N)$ by $V$. The resulting unitary, $U^{\prime}\coloneqq UV$, lies in the same equivalence class as $U$:
@@
Equation (21) defines a quotient from the unitary group to the complex projective space $\mathbb{CP}^{N-1}$:
@@
We can now induce a norm on the tangent to the state space that accounts for this equivalence.
Suppose we have a state $|\psi(t)\rangle$, and we change it slightly with derivative $|\dot{\psi}(t)\rangle=-iH(t)|\psi(t)\rangle=-i(\sum_{I}Y^{I}(t)T_{I})|\psi(t)\rangle$. Certain linear combinations of $Y^{I}$ are fixed by fixing the state and its derivative, while others remain free; the latter correspond to the stabilizer of $\lvert\psi(t)\rangle$. Minimizing the cost function over the stabilizer’s degrees of freedom produces a norm on the tangent to the state space:
$F^{\rm state}[|\dot{\psi}(t)\rangle]_{\lvert\psi(t)\rangle}\coloneqq\min_{\mathrm{stab}\,\lvert\psi(t)\rangle}F[Y^{I}(t)]\,.$ (23)
@@
$\mathcal{C}^{\rm state}_{F}[\lvert\psi_{\rm T}\rangle,\lvert\psi_{\rm R}\rangle]=\min_{{\{|\psi(t)\rangle\,:\,|\psi(0)\rangle=\lvert\psi_{\rm R}\rangle,\,|\psi(1)\rangle=\lvert\psi_{\rm T}\rangle\}}}\int_{0}^{1}dt\;F^{\rm state}[|\dot{\psi}(t)\rangle]_{\lvert\psi(t)\rangle}\,.$ (24)
We can use this formula to study geodesics in the state space without referring to unitaries.
@@ 5.2 Geometric features of Nielsen complexity
The penalty factors $q_{I}\geq 0$ are in one-to-one correspondence with the generators $T_{I}$ of the $\mathfrak{su}(N)$ algebra. They quantify the difficulty or cost of moving along each direction of the $\mathrm{SU}(N)$ tangent space. If we think of the tangent directions as operations performed on states, we can choose penalty factors to reflect the difficulty of implementing the corresponding gates. For instance, few-qubit operators are naturally easier to implement than many-qubit operators. Therefore, it is natural to assign smaller penalty factors to the easy generators (that implement few-qubit operations) and larger penalties to the hard generators that act on several qubits. The terms easy and hard refer to the magnitude of the penalty factor associated with each generator. A trajectory in the group manifold is then approximated by a circuit whose gates act for short time intervals.
@@
Returning to the cost function (25), consider taking $p=1$. When $p=1$, $F_{p{=}1,\vec{q}}[Y^{I}]$ has a natural physical interpretation: it counts (with an appropriate measure) the number of gates used to build the continuous trajectory in the group manifold SU($N$). However, $F_{p{=}1,\vec{q}}[Y^{I}]$ has the disadvantage of not being smooth. We therefore cannot apply the calculus of variations to the geodesics. When $p=2$, the norm (25) induces a Riemannian metric on the group manifold, simplifying the study of the geodesics. In this case, the map (22) is a Riemannian submersion [101] (defined in [102]). Heuristically, a Riemannian submersion is a smooth mapping from a higher-dimensional manifold to a lower-dimensional one. This mapping preserves distances in the directions perpendicular to the fibers. [^10] The fibers are points in the higher-dimensional space that are mapped to the same point in the lower-dimensional manifold. In our setting, the fibers correspond to the maximal subgroup $\mathrm{SU}(N-1)\times\mathrm{U}(1)$. The submersion maps $\mathrm{SU}(N)$ to $\mathbb{CP}^{N-1}$. These facts provide a systematic means of determining the metric induced by the cost-function minimization (23) on the state space. The Nielsen complexity literature has mainly focused on $p=1,2$ [99, 100, 103, 104, 105, 106, 107, 108, 109, 110, 55, 40, 111, 112, 101, 42, 113, 97, 114, 19, 44, 115].
@@
One may wonder whether Nielsen’s complexity is related to discrete gate complexity. The answer is yes: Nielsen’s complexity provides both upper and lower bounds on the gate complexity, provided the cost function satisfies certain conditions [93, 94]. The authors of [93, 94] further show that, by sufficiently penalizing the generators acting on several qubits, one can force the geodesic trajectory that reaches a unitary in the complexity geometry with arbitrary precision to only use generators acting on one or two qubits. The penalties required for this construction typically make the manifold highly curved. However, as we will discuss, Nielsen’s complexity is universal at long distances within a certain equivalence class of metrics, and this equivalence class can be shown to also contain moderately-curved metrics. A priori, this fact makes the computation of Nielsen’s complexity easier by using tools from differential geometry. Finally, let us mention that when the motion in the unitary manifold is restricted to a small subgroup, exact or semi-analytic results can be obtained, which we discuss in later subsections.
@@
There is substantial freedom in choosing a cost function. However, there is a large universality class of cost functions leading to metrics on the group manifold that may vary significantly at short distances, but differ at large distances only polynomially in the path length and in the physical system’s size [97, 19]. To simplify the rest of this subsection, we will take the penalty factors to only depend on the number of qubits $k$ on which a gate acts, and denote them as $q_{k}$. Sometimes $k$ is referred to as the weight of the gate – $q_{k}$ should not be confused with the $q_{I}$ introduced in Eq. (25). The notation $q_{k}$ means that generators $T_{I}$ that have the same weight $k$ share the same penalty factor.
Consider increasing any penalty factor while keeping the others constant. Above a critical value $\bar{q}_{k}$, the complexity ceases to depend on that penalty factor because we can implement this generator with weight $k$ as a composition of other, cheaper gates. The reason is as follows. Consider any fixed $m\geq 2$. Using the $m$-local gates, one can reconstruct all the $m^{\prime}$-local gates for every $m^{\prime}$ [15]. Consequently, one can identify a certain critical schedule $\{\bar{q}_{k}\}$ in terms of which we can define a universality class of cost functions that yield metrics with the same long distance properties [97]. Specifically, given the critical schedule, we consider penalty factors $\{q_{k}\}$ satisfying $q_{2}=\bar{q}_{2}$ and, for all $m>2$, $q_{m}\geq\bar{q}_{m}$. Metrics arising from this universality class of cost functions may differ greatly at short separations but lead to approximately equal distances at long separations. That is, short and long distance scales decouple. The critical schedule is conjectured to be the only universality class member for which the conjugate points are pushed to infinity, and geodesics are straight lines for an exponential time in the number of qubits [97]. [^11] A conjugate point occurs along a geodesic if there is a local shortcut from the start of the trajectory to some point along it (the notion of local shortcut will be more precisely defined below). Such conjugate points can be diagnosed by locally perturbing a geodesic. If the perturbation does not change the length, the geodesic can be deformed to find a shorter geodesic between the same endpoints. There can also be global obstructions to minimality of a geodesic, such as geodesics encircling a sphere in opposite directions along a diameter. We will discuss conjugate points in more detail in Sec. 5.4. Brown argued that the same universality class consists of metrics whose associated Nielsen complexity, measured by the geodesic length from the origin, is polynomially equivalent to the gate complexity. At the same time, it is possible to pick a metric inside the universality class such that the unitary manifold has modest curvature [19], thus making computations easier.